OpenStudy (anonymous):

Find the area of the region enclosed by the curves y = sqrt(x) & y = 10 - sqrt(x), on the y-axis.

OpenStudy (anonymous):

I know to make the functions in terms of x. So \[x = y^2\] & \[x = (10-y)^2\] But I don't know what I am doing wrong for the final answer... I either get 2000/3 or 0. None are right on our online math site.